THE CALDERÓN COMMUTATOR ALONG A PARABOLA

ANTHONY CARBERY STEVE HOFMANN JAMES WRIGHT

Abstract. We introduce an analogue of Calderón’s first commutator along a parabola, and establish its L^{2} boundedness under essentially sharp hypotheses.